Use the Laws of Logarithms to combine the expression. 3 In(2) + 4 1n(x) - In(x + 4) 2
Answer #1

We know that log (mn) = n log m , log (mn) = log(m)+log(n) and log (m/n) = log(m)- log(n).

Therefore, 3 ln (2) + 4 ln (x) –(1/2) ln (x+4) = ln(23)+ln(x4) – ln (x+4)1/2 = ln8 + ln(x4) – ln (x+4)1/2 = ln [8x4 / (x+4)1/2].

The answer is ln [8x4 / (x+4)1/2].